OpenStudy (anonymous):

(x-21/7)+(x+4/4)=x-4

OpenStudy (turingtest):

\[{x-21\over7}+{x+4\over4}=x-4\]correct?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

find the common denominator between the 3 fractions

OpenStudy (anonymous):

i need help though

OpenStudy (turingtest):

do you know how to find the Least common multiple (LCM) of a set of numbers?

OpenStudy (anonymous):

i found the common denominator

OpenStudy (turingtest):

what is it?

OpenStudy (anonymous):

28

OpenStudy (turingtest):

right :) now multiply everything (the *entire* equation) by that

OpenStudy (anonymous):

why do i multiply that to everything? i thought its supposed to be the denominator

OpenStudy (turingtest):

by finding the LCM of the denominators and multiplying the whole equation by that, you will find that all the fractions will disappear it works like magic every time :)

OpenStudy (anonymous):

ohhh okay thanks ;)

OpenStudy (turingtest):

think of the LCM as \[28=7\cdot4=\frac{7\cdot4}1\]we multiply and things cancel\[\frac{\cancel 7\cdot4}1\frac{x-21}{\cancel 7}+\frac{7\cdot\cancel4}1\frac{x+4}{\cancel4}=\frac{7\cdot4}1(x-4)\]

OpenStudy (turingtest):

what is left is\[4(x-21)+7(x+4)=28(x-4)\]think you get the idea?

OpenStudy (anonymous):

indeed ;)

OpenStudy (turingtest):

I'm glad :)

OpenStudy (anonymous):

me too ... glad we met :)

OpenStudy (turingtest):

I'm glad you found OpenStudy hope to see ya 'round !

OpenStudy (anonymous):

care to help me with another please?

OpenStudy (turingtest):

I have to go for a moment, sorry also, please post each Q separately thanks